OpenStudy (anonymous):

is (sin(2x))^2 = sin(4x) ?

Directrix (directrix):

@imabob Welcome to Open Study is (sin(2x))^2 = sin(4x) ------------------ No. (sin(2x))^2 = (sin (2x)) * (sin (2x)) = sin ² (2x) <---- sin(4x) = 2* cos(2x)*sin(2x) <--- which can be further simplified by a second applicaton of the double angle formula for sine.